a parallel-plate has a capacitance of 25.0 uF. a nonconducting liquid with dielectric constant 6.50 is poured into the space between the plates, filling up a afractiom f of its volume a) Find the new capacitance as a function of f. b) when should you expect the capacitance to be when f=0? does your expression from part? a) agree with your answer? c) what capacitance should you expect when f=1? does the expression from part a) agree with your answer?
1
Expert's answer
2020-04-03T13:27:28-0400

(a)

"\u0421=\u0421_1+\u0421_2"


"С_1=\frac{\epsilon \epsilon_0Sf}{d}"


"С_2=\frac{\epsilon_0S(1-f)}{d}"


"C=\frac{\epsilon\epsilon_0Sf}{d}+\frac{\epsilon_0S(1-f)}{d}=\frac{\epsilon_0S}{d}(1+f(\epsilon-1))"


(b)


if "f=0"


"C=\frac{\epsilon_0S}{d}(1+f(\epsilon-1))=\frac{\epsilon_0S}{d}"


(c)


if "f=1"


"C=\frac{\epsilon_0S}{d}(1+f(\epsilon-1))=\frac{\epsilon\epsilon_0S}{d}"